OpenStudy (anonymous):

the quantity 9+ radical 3 quantity squared.

OpenStudy (anonymous):

\[{(9 + \sqrt{3})^2}\]

OpenStudy (anonymous):

yah thats it

OpenStudy (anonymous):

\[{84 + 18\sqrt{3}}\]

OpenStudy (anonymous):

\[(a + b)^2 = a^2 + 2ab + b^2\]

OpenStudy (anonymous):

thanks guys :)